OpenStudy (anonymous):

giving medal Solve 7(x+2)-10=7x+4 a. x=-2 b. x=1 c. all real numbers d. no solution

OpenStudy (chmvijay):

d

OpenStudy (mathstudent55):

Distributer the 7 on the left side. Then combine like terms. What do you get?

OpenStudy (anonymous):

hope this helps :) http://www.geteasysolution.com/7(x+2)-10=7x+4

OpenStudy (anonymous):

7x gets cancel .... no solution

OpenStudy (mathstudent55):

@shkrina @chmvijay @thmdkhs 7(x+2)-10=7x+4 7x + 14 - 10 = 7x + 4 7x + 4 = 7x + 4 4 = 4 Since 4 = 4 is a true statemnt, that means this is true for all numbers. The answer is c.